Martin’s average (arithmetic mean) score after 4 tests is 89. What score on the 5th test would bring Martin’s average up to exactly 90?
(A) 90 (B) 91 (C) 92 (D) 93 (E) 94

Answer by mananth(16946)

average score of 4 tests =89
total in these 4 tests = 4*89=356
..
average of 5 tests should be 90
so total of 5 tests should be 5*90= 450
..
450-356 = 94
He should score 94 in the 5th test